### v1

# Masked LoGoNet: Fast and Accurate 3D Image Analysis for Medical Domain

Abstract.
Standard modern machine-learning-based imaging methods have faced challenges in medical applications due to the high cost of dataset construction and, thereby, the limited labeled training data available. Additionally, upon deployment, these methods are usually used to process a large volume of data on a daily basis, imposing a high maintenance cost on medical facilities. In this paper, we introduce a new neural network architecture, termed LoGoNet, with a tailored self-supervised learning (SSL) method to mitigate such challenges. LoGoNet integrates a novel feature extractor within a U-shaped architecture, leveraging Large Kernel Attention (LKA) and a dual encoding strategy to capture both long-range and short-range feature dependencies adeptly. This is in contrast to existing methods that rely on increasing network capacity to enhance feature extraction. This combination of novel techniques in our model is especially beneficial in medical image segmentation, given the difficulty of learning intricate and often irregular body organ shapes, such as the spleen. Complementary, we propose a novel SSL method tailored for 3D images to compensate for the lack of large labeled datasets. The method combines masking and contrastive learning techniques within a multi-task learning framework and is compatible with both Vision Transformer (ViT) and CNN-based models. We demonstrate the efficacy of our methods in numerous tasks across two standard datasets (i.e., BTCV and MSD). Benchmark comparisons with eight state-of-the-art models highlight LoGoNet’s superior performance in both inference time and accuracy. https://github.com/aminK8/Masked-LoGoNet.
1. Introduction
Accurate medical image segmentation can facilitate disease diagnosis and treatment planning (Yun et al., 2021; Hatamizadeh et al., 2021; Zhou et al., 2022). One of the fundamental difficulties in this task is the presence of organs or structures that span a large receptive field. These structures may have irregular shapes, complex boundaries, or significant variations in appearance, making the segmentation task particularly demanding. Additionally, the high cost of expert annotation in this domain restricts the availability of large-scale labeled datasets. Consequently, it limits the applicability of general domain computer vision methods (Xu, 2021; Haghighi et al., 2022; Cai et al., 2023). Furthermore, deployed systems usually process a large volume of images on a daily basis, which demands a substantial computational resources and leaves a large carbon footprint (Henschel et al., 2020; Chen et al., 2019; Li et al., 2021b). In the present work, we propose a fast and accurate image segmentation architecture for the medical domain. We also propose a pre-training algorithm to exploit unlabeled images, and therefore, alleviate the demand for human annotation.
Our architecture is based on the widely adopted U-shaped model. We particularly employ two strategies to enhance the inference speed, and simultaneously, maintain the prediction accuracy. First, in contrast to existing models that rely on Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) as encoders (Dosovitskiy et al., 2020; Isensee et al., 2018), we employ the large-kernel attention model (LKA) (Guo et al., 2022) in our feature extractor, which we term ULKANet (Unet Large Kernel Attention Network). As we discuss in the next section, CNN and ViTs-based models suffer from a high memory complexity, are slower during inference, and lack a proper strategy to process image sequences.111The term ”sequence” in 3D medical imaging refers to a series of volumetric data that can be either a temporal sequence, capturing changes over time in a specific anatomical region, or a spatial sequence, consisting of different slices from a 3D volume to provide a comprehensive view of the anatomy from various angles. On the other hand, our method is demonstrably more efficient due to the presence of LKA in the encoder.
Our second strategy is to enhance feature extraction through an inductive bias. Learning short-range and long-range dependencies is essential in medical image segmentation due to the large receptive field of organs. Existing studies employ U-Net with the attention mechanism, and vertically scale up their architecture to increase the network capacity for handling feature dependencies (Oktay et al., 2018; Sinha and Dolz, 2020; Cai and Wang, 2022). In contrast to these methods, we incorporate our encoder (ULKANet) into a dual encoding algorithm to learn local (short-range) as well as global (long-range) features. This enables us to keep the network size manageable, and at the same time, maintain the prediction accuracy. We term this model LoGoNet (Local and Global Network). Our model is particularly advantageous for segmenting organs such as the spleen, which has an elongated shape and irregular corners. Such body organs demand the extraction of global and local features for segmentation.
Finally, we propose a novel self-supervision technique for 3D images to address the lack of labeled training data. Our self-supervision method combines masking and multi-task learning. Using a multi-clustering algorithm, we generate a list of pseudo-labels for each unlabeled image. We then methodically mask selected parts of these images to implicitly feed the structural information of the unlabeled data into our model. An property of our proposed SSL technique lies in its versatility, as it seamlessly supports both CNN and ViT-based models. This flexibility sets our strategy apart from conventional SSL approaches, which often cater to a specific architecture (He et al., 2022; Zhou et al., 2021; Kakogeorgiou et al., 2022). Furthermore, our strategy leverages the inherent characteristics of 3D medical images, specifically embracing the concept of sequential images and neighborhood information of voxels in 3D images.
We evaluate our techniques on numerous tasks across two datasets, i.e., the BTCV dataset (Gibson et al., 2018) for segmenting body organs, and the MSD dataset (Simpson et al., 2019) that encompasses diverse tasks in medical imaging, ranging from liver tumors to cardiac and lung segmentation. Additionally, we benchmark our method against eight state-of-the-art baseline models. The results demonstrate the effectiveness and efficiency of our techniques. To offer a thorough insight into the unique attributes of our approach, we undertook extensive experiments, meticulously showcasing our model’s distinguishing features and capabilities. To summarize, our contributions are threefold:
We propose a resource-efficient model based on the commonly used U-shaped architecture. Our model has a short inference time and, at the same time, outperforms state-of-the-art methods. We achieve this by employing two strategies: first, instead of relying on CNN or ViT-based techniques, we utilize the large-kernel attention method to reduce computational complexity. Second, instead of vertically scaling up our network to improve feature extraction, we use a dual encoding algorithm to facilitate the task. We empirically demonstrate that our strategies combined achieve the best inference time and the highest precision.
We propose a multi-task self-supervision technique to exploit unlabeled images, and to overcome the lack of labeled data by employing a new masking approach specifically designed for 3D images.
We evaluate the efficacy of our model on numerous tasks across two datasets, and show that it outperforms eight state-of-the-art baseline models.
2. Related Work
To model long-range dependencies in images, existing studies mostly use vision transformers (Cao et al., 2023; Hatamizadeh et al., 2021; Valanarasu et al., 2021; Han et al., 2021; Karimi Monsefi et al., 2023; Wu et al., 2022; He et al., 2021b; Eskandari et al., 2023; Azad et al., 2023), and draw ideas from sequence modeling in Natural Language Processing (NLP). A limitation of these approaches is their treatment of images as 1D sequences, thereby overlooking the input’s inherent 2D or 3D structure. They struggle to grasp the spatial relationships between pixels, leading to poor performance in tumor detection or organ segmentation tasks. Additionally, they suffer from quadratic memory complexity, leading to high processing costs and slowness for high-resolution images, especially in the 3D context (Singh et al., 2020; Zhang et al., 2021; Khan et al., 2022; Jangid et al., 2024; Wang et al., 2023). In contrast, our proposed model, ULKANet, adopts an attention mechanism with LKA222LKA (Guo et al., 2022) is a method for computer vision tasks that effectively captures long-range relationships from input features. LKA reduces computational costs while generating attention maps highlighting essential features without additional normalization functions by decomposing large kernel convolutions into spatial local, long-range, and channel convolutions. to handle long-range dependencies while preserving the spatial structure of the images. This distinctive property enables our model to capture spatial patterns of the input more effectively, resulting in more informative representations. This is particularly advantageous in detecting tumors, where the conditions may extend over a considerable area, and models that rely solely on local features often fail to detect such cases (Wang et al., 2022).
In addressing dependencies within data, various techniques are employed based on the range of the dependencies. CNN-based models have proven effective for short-range dependencies, leveraging convolutional operations to identify relevant spatial patterns efficiently. Through this approach, hierarchical representations are learned, enhancing the understanding of the intrinsic structure of the data (Xia et al., 2020; Myronenko, 2019; Isensee et al., 2018; Zhou and Ramnath, 2022). However, our methodology takes a comprehensive approach, recognizing the importance of long and short-range dependencies. We adopt a dual encoding strategy to achieve this, incorporating an attention mechanism in parallel mode. This dual encoding technique enables the simultaneous capture and encoding of both types of dependencies, providing a more holistic representation of the underlying relationships in the data.
Next, the lack of labeled training data is a primary challenge in medical image analysis. To address this challenge, some studies have focused on domain-specific pretext tasks, as seen in Zhu et al. (2020); Ahn et al. (2021); Zhao et al. (2021); Cao et al. (2020); He et al. (2021a); Xu and Adalsteinsson (2021). Others, such as Zhou et al. (2020), adapt contrastive learning techniques to suit medical data by focusing on feature level contrast, creating homogeneous and heterogeneous data pairs by mixing image and feature batches, and utilizing a momentum-based teacher-student architecture. A comprehensive evaluation of various SSL strategies for 3D medical imaging was conducted by Taleb et al. (2020). Azizi et al. (2021) demonstrated the benefits of pre-training a model on ImageNet for dermatology image classification, showcasing the potential of transfer learning in the medical imaging domain.
Tang et al. (2022) combined inpainting with contrastive learning to improve medical segmentation. Recently, Chen et al. (2023) introduced a masked approach as a pretext task for 3D medical images. Their method centers around the task of reconstructing the masked regions of images, essentially treating it as a single-task objective. However, in contrast to this approach, our work builds upon the concept of masking but focuses on predicting (pseudo-)labels for masked images. This constitutes a multi-label learning framework where each masked image is associated with more than one label.
Our work notes that capturing information in 3D medical images not only relies on individual images but also involves considering the sequence of images. To utilize this additional source of information, we propose a multi-learning approach employing a batch of clustering algorithms. These algorithms aid in establishing multiple labels for each image, enabling the model to learn the data characteristics from various aspects.
3. Proposed Model
Figure 0(a) illustrates the architecture of our model LoGoNet. The forward pass begins by processing the input data in parallel. We have two modules in this stage, the global and the local modules. In the global module, the original data cube333”Cube” typically refers to a three-dimensional (3D) region of interest (ROI) within the volumetric medical image. Medical images, such as those obtained from MRI or CT scans, are often represented as 3D volumes, where each voxel (3D pixel) contains intensity or other information about the tissue or structures being imaged. A cube in this scenario is a 3D subset of the entire image volume. is fed into our feature extractor (ULKANet). In the local module, the same data cube is partitioned into smaller cubes, and then, each cube is processed by a separate feature extractor. Afterwards, the resulting feature tensors are concatenated to reconstruct the input. Then the outputs of the global and the local modules are aggregated by an element-wise summation operator–note that they have the same dimensions. Finally, the resulting tensor is passed through a convolution kernel followed by a 3D batch normalization operator and a GELU activation function to shape the input to our final classifier. Our final classifier is a convolution kernel.
In the next section, we discuss our encoder-decoder architecture (ULKANet), which is armed with a adaptation of LKA in the encoding phase. We then explain our local-global dual encoding strategy, which enables our model to extract feature dependencies at varying scales. After describing our model in detail in Sections 3.1 and 3.2, we then explain our novel pre-training method in Section 3.3. We use this pre-training algorithm to initialize the parameters of our model before beginning to fine-tune the network on labeled data.
3.1. LKA in Feature Extractor: An Alternative to CNN and ViTs-based Models
Figure 0(b) illustrates an overview of our feature extractor (ULKANet), which is a U-shaped model and has an encoder and a decoder. The encoder consists of a sequence of blocks. Each block consists of a repeating sequence of three components: a patch embedding component, a chain of transformer-like modules that employ LKA ( modules for block of the encoder), and a layer normalization component. For conciseness, Figure 0(b) only shows the top-level blocks, while a detailed illustration of the model architecture and inner components is provided in the appendix section 7.
The Patch Embedding component plays a crucial role in the processing of input data within the encoder block, transforming the input into a tensor that is subsequently passed to the next component in the sequence. Throughout the current encoder block, the dimension of the embedding vectors remains constant, denoted as .
The mathematical representation of the projection operation is defined as follows:
where represents the input with five dimensions , and is the batch size, is the channel size, is the size of the 3D convolution kernel, is the number of channels for the output of Conv3D, and Norm represents the batch normalization operator. denotes the size of the 3D input, and the flatten operation results in a tensor with dimensions . The Patch Embedding process serves to efficiently capture and represent the relevant features of the input data, facilitating the subsequent stages of the network architecture.
To enable our model to efficiently extract complex feature dependencies that are often present in medical images, we opt for using transformer modules. However, instead of using the regular transformers with self-attention that is slow and needs more memory (Singh et al., 2020), we use LKA (Guo et al., 2022) in the attention layer. This type of attention mechanism decomposes large convolution kernels into spatial dependencies and channel convolutions. It enables our model to go deeper and remain memory efficient. The attention module is implemented as follows:
where is the input tensor and ChConv3D is a depth-wise convolution operating on a single channel. DiConv3D is a dilated depth-wise convolution to broaden the receptive field and to enable the extraction of long-range dependencies. The point-wise convolution is applied to aggregate the information across the channels. The final activations are obtained as follows:
where is the element-wise product. The remaining components of the transformer block follow the conventional structure of typical transformers.
The decoder in our model aims to restore the spatial resolution of the input using a sequence of blocks (green circles in Figure 0(b)). Each decoder block consists of a chain of three convolution modules followed by an upsampling operation. The convolution modules are responsible for volumetric convolution operation. They consist of a Conv3D layer and a batch normalization layer, followed by a LeakyReLU activation function. The upsampling operation scales the resolution by a factor of two. As we stated earlier, a second larger illustration of our architecture that shows the inner modules can be found in the appendix section 7.
For each individual block in the encoder, the decoder has one corresponding block. There is also an additional decoder block in the bottleneck layer, as shown in Figure 0(b). The input to each decoder block is supplied by the block in the previous layer and also the corresponding encoder block through a skip connection. In order to enhance the reconstruction of input, we use the skip connections to facilitate the transfer of high-level features (Yosinski et al., 2014) to the layers that are responsible for the reconstruction task.
3.2. Dual Encoding Strategy: An Alternative to Increasing Model Capacity
One of the difficulties in medical image segmentation is the presence of organs that have complex shapes. For instance, the human gallbladder and spleen have an elongated structure. Hence, to achieve satisfactory performance in the segmentation task, the model should be able to detect and extract relevant features in multiple regions of the input images, heavily relying on global features. On the other hand, this organ has irregular corners. This characteristic requires the model to be able to detect local features in multiple regions of the input. While increasing the model capacity by adding more layers, and also composing larger training sets, will potentially enable the model to automatically learn these regularities, this will likely increase costs during both the deployment and development stages.
To reduce the burden of automatic feature mining and, consequently, to reduce the costs, we propose to impose an inductive bias (Mitchell, 1997) on the feature extraction process. We propose to have two feature extractors in parallel, one focusing on the global scale and another one focusing on the local scale–as shown in Figure 0(a). The global module is able to extract long-range dependencies due to access to the original data cube. On the other hand, the local module focuses on short-range dependencies. This is accomplished by partitioning the input cube into smaller ones, allowing for a more focused analysis and resulting in finer-grained features.
To implement our idea, we use one instantiation of ULKANet in the global module, and a sequence of instantiations of ULKANet in the local module. In the analysis section, we show that while using only one ULKANet can reduce the model size and speed up inference, it will also significantly deteriorate prediction accuracy. Additionally, we show that alternative strategies, used in comparable models, are either slower or achieve lower prediction accuracy. To prepare the data for the local module, the input 3D image is split into smaller cubes of size . Given an image of size , the value of is obtained by .
To reconstruct the input data cube, the outputs of the local module are concatenated, as shown in Figure 0(a). In order to aggregate the outputs of the global and local modules, we use an element-wise summation operator. The resulting tensor is expected to represent both global and local range dependencies.
3.3. Pre-Training Method: Exploiting Unlabeled Images
Before fine-tuning our model on labeled data, we utilize a multi-task pre-training technique to put the model weights in a favorable state. This self-supervised approach allows the model to learn general information from 3D medical images, without the necessity of ground-truth labels.
Pre-training of our model is done in three stages. First, we methodically mask certain regions of the input images. In this stage, the goal is to capture long-range and short-range feature dependencies. Second, we generate pseudo-labels for the masked images. The model later learns to generalize to unseen cases by predicting the pseudo-labels of the masked data. Finally, the masked images, along with their pseudo-labels, are used to pre-train the model. Below we explain each step.
3.3.1. Masking Algorithm
In 3D imaging, objects are depicted across multiple 2D surfaces. Therefore, we argue that an effective masking strategy should step beyond 2D inputs.
In order to help the model explore not only the dependencies between pixels in 2D images but also the connections among pixels that form 3D masses, we propose an algorithm to mask chains of patches in an image sequence.444Note that in speech processing, where data is naturally sequential, applying this technique seems to be the default method (Hsu et al., 2021). However, to our knowledge, we are the first to propose this technique in the computer vision domain. We begin by randomly selecting an image from the set of unlabeled data, with probability for selecting an individual image. Along the selected image, we also retrieve the preceding images in the same sequence. Then, we apply a masking technique to the images in the chain. Various masking techniques can be used in this stage (Shi et al., 2022; Li et al., 2021a); we employ the method introduced by Xie et al. (2022). Therefore, for each image in the chain, we randomly select a patch size , and partition it into patches, where and are the height and width of the image. Finally, with the probability we mask out each patch of the image. Appendix section 8.1 shows the details about the masking algorithm, and how we tune the hyperparameters.
In contrast to algorithms such as SimMIM (Xie et al., 2022), our proposed approach distinguishes itself by selecting a sequence of images and subsequently applying masking to that sequence. This method facilitates the encoder in gathering information by focusing on the interdependence of voxels within the sequence of images. Notably, our algorithm operates independently of the specific model structure, diverging from approaches seen in studies by Kakogeorgiou et al. (2022), He et al. (2022), and Zhou et al. (2021), all of which exhibit a reliance on model structure. Furthermore, our approach is compatible with Vision Transformer (ViT)-based (Dosovitskiy et al., 2020) and CNN-Based models.
3.3.2. Pseudo-Label Generation
Our pseudo-label generation algorithm assigns labels to all the images in the unlabeled set. Later in the pre-training pipeline, our model is asked to predict the pseudo-labels of the masked out images in each sequence. The information conveyed by the distorted images is insufficient for label prediction. Therefore, the model must explore the associations between pixels across multiple 2D images in the sequence to correctly predict the pseudo-labels of the target images. In the analysis section, we empirically show that this exploration task helps the model to learn the properties of the domain and to generalize better.
A clustering algorithm is employed for the pseudo-label generation. For simplicity, we use the k-means clustering method, although other types of clustering methods, such as hierarchical or spectral methods, can be utilized. Given a random number as the predefined number of clusters, we train a k-means clusterer on a random subset (e.g. in our experiments) of the unlabeled data. Then we use the clusterer to label the entire unlabeled set. Note that masking is not applied in any of these stages, and the clusterer has access to the unmasked images. The obtained labels are used as pseudo-labels to pre-train the model by predicting the corresponding labels for every masked image.
The k-means clusterer is able to use all the properties of the images to form the clusters. For instance, a cluster may constitute images that illustrate elongated organs, while another cluster may constitute images that depict organs that have particular corners. During pre-training, the model is asked to recover the pseudo-labels of a sequence of images that are distorted by masking. In order to predict their correct labels, the model must discover the associations between neighboring pixels. This pretext task enables the model to learn long-range and short-range spatial dependencies effectively.
Assuming that a clustering method exploits a finite set of characteristics in data to form the clusters, our model needs to learn these characteristics to correctly assign each image to the associated clusters. We conjecture that having different clusterers labeling the data and then using our model to simultaneously predict these multiple labels can further help the model gain broader knowledge from the data. From a different perspective, we can assume that recovering the characteristics of each clusterer is a separate pre-training task, and then, concurrently recovering the characteristics of multiple clusterers is a multi-task training. The efficacy of multi-tasking is well-documented in the machine learning literature (Caruana, 1997). Figure 2 shows our pre-training pipeline. In this figure, denotes the total number of clusterers, and denotes the pseudo-label generated by -th clusterer for the -th masked image in the sequence.
3.3.3. Pre-Training Loss Function
To pre-train our model, we use a cumulative negative log-likelihood function on the model predictions for the masked images as follows:
where is the number of clusterers, is the number of masked images that can be calculated by , where is the length of image sequence for masking, and is the number of concurrent masked sequences, if present. are masked images, and is the probability that the -th masked image in the sequence (i.e., ) is correctly assigned to the pseudo-label generated by the -th clusterer. The value of is calculated by a softmax function on top of the pre-training classification head, which is a simple feed-forward network.555Replacing the pre-training head with a finetuning head is an established practice in the self-supervision literature (Devlin et al., 2019). Therefore, given a clusterer, we have:
where is the exponential function, is the number of clusters generated by the clusterer, is the cluster that the input image belongs to, and is the -th logit of the pre-training classification head. The hyper-parameter is called the softmax temperature. The value of determines the strength of the gradients backpropagated through the network. Lower temperature values increase the magnitude of gradients (Hinton et al., 2015). This, in turn, reduces the standard deviation of output probabilities–also known as sharpening the posterior probabilities.
Our loss function (Equation 4), iterates over all the predictions that our model makes during the pre-training stage and penalizes for the errors. As we discussed earlier, our pretraining framework enables LoGoNet to become familiar with the properties of the domain to generalize better by exploiting unlabeled data. We empirically support this argument in our analysis section. Additional experiments can be found in appendix section 8.1.
4. Experimental Setup
| Models | SegResNetVAE | SwinUNETR | UNETR | UNet++ | nnUNet |
| FLOPs (G) | 15.50 | 329.84 | 264.59 | 4229.20 | 1250.65 |
| # Param | 3.9 M | 62.2 M | 101.7 M | 84.6 M | 30.7 M |
| Models | DiNTS Search | DiNTS Instance | Attention U-Net | LoGoNet | |
| FLOPs (G) | 743.88 | 743.88 | 7984.21 | 246.96 | |
| # Param | 74.1 M | 74.1 M | 64.1 M | 67.5 M | |
In this section, we briefly describe the datasets used in the experiments, provide a list of baseline models we compare to, and also provide an overview of our setup.
Datasets. We use two widely used standard datasets. As the first dataset, we use the BTCV dataset666Available at https://www.synapse.org/#!Synapse:syn3193805/wiki/217789 introduced by Gibson et al. (2018). This dataset contains 13 segmentation tasks, and each task has 40 data points obtained via abdominal CT scans. As the second dataset, we use the MSD dataset777Available at http://medicaldecathlon.com/ introduced by Simpson et al. (2019). This dataset contains a variety of tasks obtained via magnetic resonance imaging (MRI), computed tomography (CT), and positron emission tomography (PET). We use six different tasks from this dataset that contain a total of 900 examples. As the unlabeled data, we use the meta-dataset collected by Tang et al. (2022), which consists of 4,500 examples. The images in this dataset are not annotated, and are 3D scans covering a variety of organs. Detailed information about the datasets can be found in appendix section 10.
Baselines. We compare LoGoNet to a suite of baseline models, including those that use Visual Transformers or Convolutional Neural Networks. We compare to nnUNet (Isensee et al., 2018), Attention U-Net (Oktay et al., 2018), SegResNetVAE (Myronenko, 2019), UNet++ (Zhou et al., 2019), DiNTS (two variations of Search and Instance) (He et al., 2021c), SwinUNETR (feature size 48) (Hatamizadeh et al., 2021), and UNETR (feature size 32) (Hatamizadeh et al., 2022). A brief description about each baseline model can be found in appendix 9.
Setup. We follow standard practices to carry out the experiments. We use the Dice metric, a common metric for the image segmentation task, to report the performance results. We conduct the experiments in each dataset task separately and report the average results for five runs in the BTCV dataset and two runs in the MSD dataset. Detailed information about hyperparameter tuning, configurations, and implementation can be found in appendix section 8.
Our default LoGoNet and ULKANet models have four encoder blocks with 3, 4, 6, and 3 transformer modules in each block, respectively. The dimensions of the embedding vectors in these models are 64, 128, 256, and 512, respectively.
5. Results
5.1. Main Results
| Models | Spl | RKid | Lkid | Gall | Eso | Liv | Sto | Aor | IVC | Veins | Pan | Rad | Lad | AVG |
| UNETR | .912 | .940 | .938 | .693 | .690 | .954 | .754 | .891 | .830 | .703 | .734 | .660 | .577 | .790 |
| SegResNetVAE | .941 | .938 | .933 | .670 | .718 | .955 | .745 | .892 | .848 | .695 | .783 | .633 | .528 | .791 |
| nnUNet | .859 | .944 | .924 | .796 | .755 | .960 | .781 | .894 | .849 | .756 | .776 | .675 | .663 | .818 |
| Attention U-Net | .955 | .936 | .930 | .735 | .739 | .964 | .770 | .898 | .852 | .753 | .763 | .695 | .688 | .821 |
| DiNTS Instance | .935 | .942 | .938 | .770 | .769 | .962 | .743 | .909 | .857 | .759 | .782 | .641 | .691 | .823 |
[…]
Table 1 compares our model to the baseline methods in terms of inference time (FLOPs) and the number of trainable parameters in the BTCV dataset. We see that our model has the lowest inference time after SegResNetVAE. Tables 2 and 3 compare the accuracy of our method to the baselines. We observe that the performance of SegResNetVAE is significantly lower than that of ours. Taking into account both the inference speed and the prediction accuracy, our model seamlessly ranks first among all the models. Appendix section 8 reports more experiments about the training time, test time, and memory consumption.
Table 1 shows that our model is considered an average-sized network. One noteworthy observation is that in some cases, e.g., nnUNet or DiNTS Instance, even though the number of trainable parameters is on a par or smaller than ours, their inference speed is substantially slower. Tables 2 and 3 show that our model exhibits superior performance compared to the baseline methods. Specifically, when evaluating our proposed model without pre-training, it outperforms the baselines across out of tasks. Furthermore, incorporating our pre-training strategy into LoGoNet enhances its performance even further, surpassing the baselines in out of tasks. These findings underscore the effectiveness and versatility of our approach in tackling a diverse range of tasks with notable efficacy.
| Models | Col | Spl | Hep | Pan | Lun | Car | AVG |
| UNETR | .677 | .969 | .715 | .699 | .730 | .953 | .790 |
| SegResNetVAE | .742 | .968 | .745 | .740 | .765 | .951 | .818 |
[…]
| Attention U-Net | - | - | - | - | - | - | - |
[…]
| SwinUNETR | .695 | .967 | .737 | .738 | .763 | .957 | .810 |
| DiNTS Search | .776 | .980 | .749 | .749 | .768 | .960 | .830 |
[…]
| LoGoNet + PRE | .801 | .980 | .779 | .833 | .828 | .958 | .863 |
In the BTCV dataset, LoGoNet outperforms the top three baseline models on average by , , and , respectively. Regarding the inference time, our model outperforms the top three models by , , and , respectively.
5.2. Analysis
In this section, we demonstrate the properties of our model from multiple aspects. Specifically, we report a qualitative comparison between our model and the best baseline model, evaluate our strategy for extracting local and global features, evaluate our pre-training approach, show the impact of model size on performance, analyze the hyper-parameter sensitivity of our model, and finally, report an ablation study on the steps in our pre-training method. The experiments in this section are carried out in the BTCV dataset unless stated otherwise.
| Ground Truth | LoGoNet | DiNTS Search |
We begin by qualitatively inspecting our model. Figure 3 compares the output of LoGoNet to the best performing baseline model in BTCV dataset, i.e., DiNTS Search (more qualitative comparisons can be found in appendix section 11). We see that our model particularly excels in segmenting organ boundaries. This can be attributed to our effective strategy for extracting local-range dependencies, which plays a crucial role in extracting details from input data. Our model’s adeptness in capturing long-range dependencies allows it to grasp contextual information that extends over significant distances within the data. Simultaneously, its proficiency in handling short-range dependencies ensures precision in capturing localized patterns.
| Models | Gall | Eso | Veins | Lad | AVG |
| ULKANet | .761 | .782 | .690 | .684 | .824 |
| LoGoNet | .818 | .786 | .769 | .698 | .854 |
To further quantitatively support our strategy for extracting local and global features in parallel, in the next experiment, we report the performance of our model compared to the regular method for extracting features from medical images, which is relying on a single feature extractor. This translates into comparing LoGoNet to our feature extractor ULKANet. Table 4 reports the results. We observe that our strategy enables our model to outperform the alternative method.
In the next experiment, we report the efficacy of our pre-training method. To carry out this experiment, we use the algorithm proposed in Section 3.3 to initialize the weights of our model, and then, we follow the regular fine-tuning steps. In Tables 2 and 3 (the last rows), we report the results of this model for both datasets, indicated by postfix PRE. We see that the improvements achieved by pre-training are consistent across both datasets.
| SSL Approach | Gall | Eso | Veins | Lad | AVG |
[…]
| Rubik’s Cube (Tao et al., 2020) | .815 | .820 | .780 | .725 | .859 |
| SimCLR (Chen et al., 2020) | .829 | .803 | .780 | .720 | .859 |
| Our SSL Approach | .866 | .845 | .791 | .757 | .875 |
In the next experiment, we compare the effectiveness of our self-supervised pre-training approach to the alternative methods. In particular we compare to SimMIM (Xie et al., 2022), Rubuk’s Cube (Tao et al., 2020), and SimCLR (Chen et al., 2020) strategies. Table 5 reports the result. The numbers are obtained by initializing LoGoNet. Notably, our proposed model exhibits superior performance in three out of four experiments, showcasing its effectiveness in a diverse set of tasks. The complete results are available in the appendix 11. The comparison in Table 5 highlights the competitive edge of our model. This indicates the robustness and efficacy of our multi-task self-supervised learning methodology in capturing meaningful representations across various domains.
An inherent advantage of our pre-training approach lies in its versatility, as it is designed to be compatible with both CNN and ViT-based models. This flexibility broadens the applicability of our approach, allowing it to seamlessly integrate with different architectural paradigms commonly used in computer vision tasks.
| Models | Gall | Eso | Veins | Lad | AVG |
| LoGoNet | .818 | .786 | .769 | .698 | .854 |
| LoGoNet L | .847 | .781 | .768 | .710 | .855 |
[…]
| LoGoNet L + PRE | .921 | .859 | .805 | .784 | .891 |
To understand the impact of model size on the prediction accuracy, we report the performance of our default model compared to a larger variant. Our larger variant uses four encoder blocks with 3, 3, 24, and 3 transformer modules, respectively. The dimensions of the embedding vectors in this model are 96, 192, 384, and 768, respectively. Table 6 reports the results. Upon increasing the dimensions of our model, we observed an improvement in results, though it fell short of our initial expectations. We attribute this to the limited number of labeled data available. However, upon integrating our pre-training methodology into our standard and larger variants of LoGoNet, we noted a significant enhancement in performance, particularly noticeable in the larger LoGoNet.
| Model | N = 1 | N = 40 | N = 80 | |||
| Gall | Eso | Gall | Eso | Gall | Eso | |
| LoGoNet + PRE | .830 | .819 | .843 | .860 | .866 | .845 |
In Section 3.3, we claimed that having multiple clusterers serves as a multi-task training approach. In order to demonstrate the benefit of having multiple clusterers, and also show the sensitivity of our model to the number of these learners in our algorithm, we report the results of our model with varying numbers of clusterers in Table 7. We see that as the number of clusterers increases, the performance improves. The results support our hypothesis regarding the ability of our model to extract broader knowledge from the unlabeled data in the presence of multi-tasking.
Finally, we report an ablation study on the effectiveness of our masking approach during the pre-training stage. In Section 3.3, we argued that by distorting input images, the model must learn the properties of neighboring pixels in order to predict the correct labels. We then argued that this exploration task enables the model to faster learn the domain and to generalize better. The results reported in Table 8 supports our claim. We see that by incorporating the masking step, the performance noticeably improves signifying a better generalizablity of our method.
| Model | w/ M | wo/ M | w/ M + wo/ M | |||
| Gall | Eso | Gall | Eso | Gall | Eso | |
| LoGoNet + PRE | .866 | .845 | .845 | .802 | .851 | .820 |
In summary, we demonstrated the efficacy of our model in two datasets across 19 segmentation tasks. We also compared our method to eight recent baseline models, including those that use Visual Transformers. Our results testify to the effectiveness of our novel feature extraction techniques. Our analysis shows that our pre-training method is successfully able to exploit unlabeled data to improve parameter initialization. We also showed that our method significantly speeds up inference time compared to the best-performing models.
The computer vision domain is a rapidly evolving research field. It seems unrealistic to expect long-term plans. However, with the existing challenges in the medical domain, this community will invest more in developing methods for mitigating the lack of large labeled sets. Therefore, in the next step, we plan to explore Domain Adaptation, which is one of the well-known methods for addressing this challenge.
6. Conclusions
In this paper, we proposed a fast and accurate approach for 3D medical image segmentation termed LoGoNet, which combines global and local attention mechanisms. The localized mechanism in LoGoNet significantly improves segmentation, especially for small organ sections, while the incorporation of both global and local scales captures effective long-range dependencies. Additionally, we proposed a pre-training method to exploit unlabeled data to enhance model generalization. Experiments in the BTCV and MSD datasets demonstrated that LoGoNet surpasses the baselines, achieving superior segmentation accuracy. The combination of LoGoNet with pretraining further enhances performance. The utilization of masked data in pretraining framework significantly boosts the model performance to capture long-range dependencies, leading to a deeper understanding of structural relationships within 3D images, thereby improving segmentation accuracy.
Appendix
The following section presents a more detailed description of our feature extractor, ULKANet. Then, we provide the details of our experiments, including the configurations of the baseline models, our pre-training algorithm, and our model architecture. We continue with a description of each used dataset, and finally, we conclude the article by reporting an additional qualitative experiment.
7. Detailed Architecture of ULKANet
Figure 4 illustrates our feature extractor. This feature extractor is structured into two main components: an encoder and a decoder. The encoder is comprised of a series of blocks, each consisting of a recurring sequence of three essential elements: a patch embedding component, which you can find the algorithm of this component in the algorithm 1, a set of transformer-like modules employing the LKA technique (The number of these modules in the sequence is represented as ), and a layer normalization component. The LKA component contains two crucial parts, first attention, which we describe in part 3.1, and the MLP part, which you can find in the algorithm 3; also, the algorithm of LKA part is available in the algorithm 2. This architecture has been meticulously designed to process and extract crucial input data features effectively. The patch embedding operation transforms the input into a feature vector with a dimension of . Additionally, we incorporate a block, which encompasses three layers: a layer, batch normalization, and the activation function.
Furthermore, the presence of a decoder block denoted as in Figure 4 is a crucial element. This block consists of three blocks and an upsampling layer, which upscales the input by a factor of . This comprehensive structure enables our model to efficiently handle the input data and extract meaningful features for further processing.
8. Implementation Details
| Layer Number | 1 | 2 | 3 | 4 | ||||||||
| L | dim | mlpRatio | L | dim | mlpRatio | L | dim | mlpRatio | L | dim | mlpRatio | |
| Normal | 3 | 64 | 8 | 4 | 128 | 8 | 6 | 256 | 4 | 3 | 512 | 4 |
| Large | 3 | 96 | 8 | 3 | 192 | 8 | 24 | 384 | 4 | 3 | 768 | 4 |
Our model architecture has incorporated four encoder blocks, a feature in both the standard and the larger variants. However, it’s important to note that our model is flexible and can seamlessly adapt to the use of varying numbers of encoder layers. The primary distinction between the regular and large models lies in the number of transformer modules within each block and the dimensions of the internal embedding vectors.
To provide a comprehensive understanding, Table 9 presents a detailed comparison between our standard model and its larger counterpart. It’s noteworthy that, despite any variations, the size of the embedding vectors for each patch module and the mlpRatio remains consistent across all encoder blocks.
| Models | SegResNetVAE | SwinUNETR | UNETR | UNet++ | Attention U-Net | nnUNet |
| # Param | 3.9 M | 62.2 M | 101.7 M | 84.6 M | 64.1 M | 30.7 M |
| FLOPs (G) | 15.50 | 329.84 | 264.59 | 4229.20 | 7984.21 | 1250.65 |
| Inference Time (S) | 1.73 | 6.16 | 5.89 | 11.74 | 19.74 | 2.87 |
| Training Time (S) | 0.96 | 1.98 | 1.83 | 2.82 | 3.70 | 1.01 |
| Models | DiNTS Search | DiNTS Instance | ULKANet | LoGoNet | ULKANet L | LoGoNet L |
| # Param | 74.1 M | 74.1 M | 40.2 M | 67.5 M | 121.20 M | 172.54 M |
| FLOPs (G) | 743.88 | 743.88 | 109.94 | 246.96 | 218.70 | 487.65 |
| Inference Time (S) | 6.33 | 6.28 | 3.45 | 5.38 | 5.06 | 14.49 |
| Training Time (S) | 1.89 | 1.87 | 1.01 | 1.63 | 1.49 | 2.29 |
This structural consistency ensures that the essential characteristics of the model components are preserved, facilitating ease of integration and adaptability. Whether opting for the standard or larger version, users have the freedom to fine-tune the model’s performance by adjusting the number of encoder layers to suit their specific requirements. This flexibility is a key advantage of our model, allowing for versatility in handling diverse applications and tasks.
Table 10 presents a comparative analysis between our model and several baseline methods. The inference time reported herein reflects the average duration observed across five consecutive test runs on the BTCV dataset, while training time is computed as the average duration of the final ten epochs during model training on the same dataset. This rigorous assessment facilitates a comprehensive assessment of the performance and efficiency exhibited by each approach.
In the implementation of the local strategy within LoGoNet, a pivotal decision was made to partition each image tensor into segments. While this approach offers advantages in enhancing local processing capabilities, it concurrently introduces a significant surge in the number of trainable parameters. In addressing this challenge, a thoughtful strategy has been employed within the local section of LoGoNet.
Specifically, in the local processing segment of LoGoNet, a judicious selection has been made to utilize only two encoder blocks, in contrast to the four blocks employed in the global section, as previously mentioned. This intentional divergence in the number of encoder blocks between the local and global sections serves to strike a balance between computational complexity and model expressiveness.
By limiting the local section to two encoder blocks, we manage to mitigate the potential escalation in trainable parameters, thereby optimizing the trade-off between computational efficiency and model performance. This strategic choice is rooted in a nuanced understanding of the interplay between local and global processing within the overall architecture of LoGoNet.
In essence, our design rationale carefully tailors the number of encoder blocks in each section to the specific demands of local and global processing, ensuring a harmonious integration that optimally leverages the strengths of both approaches. This meticulous consideration of architectural choices reflects our commitment to achieving a well-balanced and efficient model in LoGoNet.
8.1. Pre-Training Details
We used the scikit-learn implementation888Available at: https://scikit-learn.org/stable/ of the Mini Batch KMeans algorithm as the clusterers in our pre-training pipeline. During the training phase of the k-means models, we adopted a transformation process that converted the input image from a format to a vector representation of dimensions , where is equivalent to . This transformation enabled us to generate a label for each cluster per image slice, resulting in a sequence of labels for a sequence of images. Subsequently, the model underwent iterations of training, with each iteration utilizing a randomly selected subset of the unlabeled data. To introduce diversity and enhance robustness, we employed a stochastic approach in determining the value of , randomly sampling from a range spanning to .
The information pertaining to pre-training is encapsulated in Table 12. To train the pre-trained model, we leveraged the optimizer and fine-tuned the process by configuring specific parameters. In particular, we assigned values of and to and respectively. Additionally, the sequence of distorted images, denoted as , was set to .
| Hyperparameter | M = 3 | M = 5 | M = 7 |
[…]
| = 0.3 | .841 | .843 | .838 |
Table 11 presents the outcome of selecting hyperparameter values, with results obtained from the BTCV dataset using the ULKANet model. This tabulated information sheds light on the meticulous decision-making process involved in determining specific values for key hyperparameters, providing valuable insights into our experimental configuration.
Our observations reveal that augmenting both the values of (length of sequenced mask images) and (rate of sampled images) results in an increased rate of masked images. However, this heightened rate poses challenges for our model, making it more intricate to exploit dependencies between successive slices for effectively capturing information related to missing voxels. This delicate interplay between hyperparameters emphasizes the necessity of finding an optimal balance to enhance model performance, as an excessive increase in masked images may impede the model’s ability to leverage contextual dependencies within the data.
Furthermore, we introduced randomness in the selection of patch sizes, choosing from the set .
Our innovative pre-training approach involves the incorporation of a header designed to adapt the model output to align with the requirements of our pseudo-labeling. Figure 2 shows the structure of our proposed pre-training. The structure of this header can be found in algorithm 4. This header modification serves as a crucial step in optimizing the model’s output for seamless integration with our pseudo-labeling methodology during the training process.
| Configuration | Value | |||||
| Optimizer | ||||||
| Epochs | ||||||
| Batch Size per GPU | ||||||
| Number of GPUs | ||||||
| Weight decay | ||||||
| Optimizer momentum | , = , | |||||
| Peak learning rate | ||||||
| Learning rate schedule | ||||||
| Warmup epochs | 10 | |||||
| Dropout | 0 | |||||
| Rand Spatial Crop Samples Data | ||||||
| MONAI Transforms: ScaleIntensityRanged |
| 0.1 | ||||||
| M (Size of masked sequence) | ||||||
| (Size of Patches) |
8.2. Configuration Setup
Table 14 provides a comprehensive overview of the specifics pertaining to our training or fine-tuning procedures. This tabulated information encapsulates crucial details, offering insights into the intricacies of our training regimen. By examining the contents of this table, readers can gain a nuanced understanding of the parameters, configurations, and methodologies employed during the training or fine-tuning phase of our experiments.
Our experimental setup involved the utilization of 8 nodes, each equipped with dual-GPU chipsets. Each GPU within these chipsets boasted a substantial 35GB of memory, ensuring ample resources for running our experiments efficiently. To adhere to established standards and foster equitable comparisons, we employed a comprehensive array of augmentation techniques to augment data variability. It’s noteworthy that these augmentations were uniformly applied to all models, encompassing both our proposed models and the baseline models. This meticulous approach ensures a fair and unbiased comparative analysis.
For the implementation of our models and the baseline models, we leveraged the framework, which provided a robust and versatile foundation for our experimentation. This framework facilitated the seamless integration of our methodologies, streamlining the implementation process and contributing to the reliability of our experimental results. 999Available at: https://monai.io/ In the course of each iteration, we strategically implemented a randomized cropping strategy, extracting two images for each case during the training phase. This deliberate approach was employed with the intent of diversifying the training dataset for each input case within every epoch, thereby enhancing the overall richness of the training process. The randomness introduced by the cropping strategy contributed to a more robust and varied training experience.
Furthermore, to fine-tune and optimize our training procedure, we incorporated the advanced optimizer. This optimizer played a pivotal role in adjusting the model’s weights during training, ensuring an efficient convergence towards optimal performance. The synergistic combination of the cropping strategy and the utilization of the optimizer exemplifies our commitment to refining the training dynamics for improved model efficacy.
To refine our model using labeled data, we employed the DiceCELoss as the objective function during the fine-tuning or training process. The DiceCELoss function serves as a crucial metric, enabling us to strike a balance between the Dice coefficient and Cross-Entropy, optimizing the model’s performance on the labeled dataset. This choice of objective function reflects our commitment to a meticulous fine-tuning or training process, ensuring that the model is adept at capturing the nuanced patterns present in the labeled data. The DiceCELoss is articulated by the following formulation:
where
and
Thus, our fine-tuning and training loss term is the weighted summation between the regular dice loss term and the cross entropy term. represents the predicted probability for the -th class. represents the ground truth label for the -th class. represents the number of classes. is a small constant (e.g., 1e-5) added to the denominator to avoid division by zero.
| LoGoNet | ||||||
| 1.0 | 1.0 | 0.0 | 1.0 | 1.0 | 0.0 | |
[…]
| Configuration | BTCV | MSD | ||||||||||
| Optimizer | AdamW | |||||||||||
| Epochs | 5000 | 1000 | ||||||||||
| Batch Size per GPU | 2 | 1 | ||||||||||
| Number of GPUs | ||||||||||||
| Weight decay | ||||||||||||
| Optimizer momentum | , = , | |||||||||||
| Peak learning rate | ||||||||||||
| Learning rate schedule | LinearWarmupCosineAnnealingLR | |||||||||||
| Warmup epochs | ||||||||||||
| Dropout | 0 | |||||||||||
| Rand Spatial Crop Samples Data | ||||||||||||
| MONAI Transforms: ScaleIntensityRanged |
| MONAI Transforms: Spacingd |
| Data Augmentation: RandFlipd | prob for each axis = 0.2 | |||||||||||
| Data Augmentation: RandRotate90d | prob = 0.2 | |||||||||||
| Data Augmentation: RandScaleIntensityd |
| Data Augmentation: RandShiftIntensityd |
Our experimentation has revealed that assigning equal weights to both CELoss and DiceLoss yields more favorable outcomes, surpassing the performance achieved with other weight ratios. The results of various weight configurations for losses are presented in Table 13. By according equal significance to both Cross-Entropy Loss (CELoss) and Dice Loss, we strike a balance that enhances the model’s ability to effectively capture diverse patterns in the data. The results highlight the critical role of assigning equal weights in attaining optimal outcomes and underscore the robustness of this approach within the framework of our experimental setup. This observation emphasizes the significance of maintaining a balanced weighting scheme, showcasing its effectiveness in ensuring stability and reliability across various experimental conditions. The consistent performance across diverse scenarios further validates the efficacy of the equitable weighting strategy adopted in our study.
| Task | Colon Cancer | Spleen | Hepatic Vessels | Pancreas Tumour | Lung Tumours | Cardiac |
| Train | 119 | 35 | 241 | 210 | 51 | 14 |
| Validation | 13 | 3 | 26 | 23 | 5 | 2 |
| Test | 14 | 9 | 61 | 57 | 13 | 4 |
| Models | Spl | RKid | Lkid | Gall | Eso | Liv | Sto | Aor | IVC | Veins | Pan | Rad | Lad | AVG |
| UNETR | .912 | .940 | .938 | .693 | .690 | .954 | .754 | .891 | .830 | .703 | .734 | .660 | .577 | .790 |
| SegResNetVAE | .941 | .938 | .933 | .670 | .718 | .955 | .745 | .892 | .848 | .695 | .783 | .633 | .528 | .791 |
| nnUNet | .859 | .944 | .924 | .796 | .755 | .960 | .781 | .894 | .849 | .756 | .776 | .675 | .663 | .818 |
| Attention U-Net | .955 | .936 | .930 | .735 | .739 | .964 | .770 | .898 | .852 | .753 | .763 | .695 | .688 | .821 |
| DiNTS Instance | .935 | .942 | .938 | .770 | .769 | .962 | .743 | .909 | .857 | .759 | .782 | .641 | .691 | .823 |
| UNet++ | .934 | .931 | .925 | .810 | .715 | .961 | .786 | .900 | .846 | .747 | .829 | .685 | .679 | .827 |
| SwinUNETR | .952 | .947 | .945 | .790 | .770 | .963 | .755 | .901 | .850 | .771 | .760 | .702 | .659 | .828 |
| DiNTS Search | .937 | .934 | .930 | .788 | .770 | .960 | .774 | .904 | .866 | .751 | .813 | .670 | .711 | .831 |
| ULKANet | .954 | .936 | .938 | .761 | .782 | .964 | .814 | .885 | .831 | .690 | .786 | .690 | .684 | .824 |
[…]
| ULKANet + PRE | .954 | .940 | .933 | .836 | .794 | .966 | .853 | .899 | .854 | .760 | .823 | .703 | .734 | .850 |
[…]
| ULKANet L | .948 | .938 | .933 | .832 | .789 | .967 | .832 | .888 | .862 | .744 | .784 | .720 | .690 | .841 |
[…]
9. Baselines
This section will comprehensively discuss our baseline models, drawing comparisons with LoGoNet across a spectrum of approaches, encompassing ViT and CNN-base models.
nnUNet (Isensee et al., 2018) is a framework based on the U-Net architecture, tailored for medical image segmentation. It simplifies the process of adapting U-Net to new tasks by focusing on crucial aspects like preprocessing, training, and inference. Through minimal modifications to U-Net and dynamic adjustments in network parameters. The nnU-Net stands out for its robustness and adaptability compared to other segmentation methods. By emphasizing essential aspects and avoiding unnecessary architectural complexities, nnU-Net consistently delivers high-quality segmentation results. While nnU-Net may provide high segmentation accuracy, its runtime efficiency might be a concern, especially in real-time or resource-constrained applications. Optimizing the model for faster inference without compromising accuracy can be challenging.
Attention U-Net (Oktay et al., 2018) proposes a novel attention gate (AG) model for medical imaging segmentation, specifically targeting structures like the pancreas with varying shapes and sizes. By integrating AGs into standard CNN architectures such as the U-Net, the model automatically learns to focus on relevant regions in input images while suppressing irrelevant areas, thereby eliminating the need for explicit external organ localization modules. Despite its strengths, the Attention U-Net may face certain limitations. While AGs offer significant improvements in prediction accuracy and computational efficiency, their effectiveness may vary depending on the complexity of the target structure and the dataset used. Additionally, the proposed model’s reliance on attention mechanisms could potentially introduce additional computational overhead during training and inference stages, particularly if not carefully optimized.
SegResNetVAE (Myronenko, 2019) employs an encoder-decoder structure augmented with a variational autoencoder (VAE) branch. The VAE component serves to regularize the shared encoder, aiding in learning from limited training data by reconstructing input images alongside segmentation during training. This regularization helps mitigate overfitting and enhances the network’s ability to generalize to unseen data, ultimately leading to improved segmentation accuracy. While the VAE regularization proves beneficial for mitigating overfitting and improving generalization, it adds complexity to the overall architecture, potentially hindering interpretability and requiring significant computational resources for training and deployment.
UNet++ (Zhou et al., 2019) introduces a novel approach to image segmentation by addressing key limitations in existing models such as U-Net and FCNs. It achieves this by employing an ensemble of U-Nets with varying depths, which share a common encoder and are trained simultaneously using deep supervision. This alleviates the need for exhaustive architecture search, allowing for flexible depth selection based on task difficulty and available labeled data. Additionally, UNet++ redesigns skip connections to enable flexible feature fusion across varying semantic scales. By allowing for the aggregation of features from different scales, UNet++ significantly enhances segmentation quality. Despite its innovative design and notable benefits, UNet++ may pose certain challenges. The increased complexity introduced by the ensemble of U-Nets and deep supervision may require additional computational resources and careful management during training. This could potentially limit accessibility for smaller research groups or organizations with limited resources.
DiNTS (He et al., 2021c) is neural architecture search (NAS) specifically tailored for 3D medical image segmentation tasks. The key innovation of DiNTS lies in its differentiable search framework, which enables efficient exploration of a highly flexible network topology search space. One notable benefit of DiNTS is its ability to support more complex network topologies, allowing for greater flexibility in adapting to the diverse characteristics of medical image segmentation tasks. Additionally, the proposed topology loss and memory budget constraints mitigate the common challenges associated with differentiable topology search, such as the discretization gap and high GPU memory usage. However, a potential weakness of DiNTS lies in its reliance on GPU resources during the search process, which could pose scalability challenges for resource-constrained environments.
UNETR (Hatamizadeh et al., 2022) introduces a novel approach to 3D medical image segmentation by integrating transformers, renowned for their effectiveness in capturing long-range dependencies, into the segmentation pipeline. Traditional convolutional neural networks (CNNs) struggle with capturing global context due to their localized receptive fields. UNETR addresses this limitation by formulating the segmentation task as a sequence-to-sequence prediction problem, where a transformer encoder is employed to learn contextual representations from 3D input volumes. This architecture follows a ”U-shaped” design, connecting the transformer encoder to a CNN-based decoder through skip connections, facilitating the computation of the final segmentation output. Transformers excel in capturing global context and long-range dependencies, which are crucial for accurate segmentation, especially in medical images with complex structures. By leveraging transformers, UNETR can effectively capture spatial relationships across the entire volume, leading to improved segmentation performance. However, UNETR also presents some weaknesses. Transformers can introduce computational complexity, potentially increasing training and inference time, which could be a limiting factor in resource-constrained environments. Moreover, transformers may struggle with capturing fine-grained local details compared to CNNs, which could impact the model’s ability to accurately segment intricate structures.
SwinUNETR (Hatamizadeh et al., 2021) introduces a novel architecture for semantic segmentation of brain tumors in MRI images, combining the hierarchical encoding capabilities of Swin transformers with a U-shaped network design and CNN-based decoders connected via skip connections at various resolutions. By leveraging the self-attention mechanism of Swin transformers, the model effectively captures long-range dependencies, enabling accurate segmentation of tumors with diverse shapes and sizes. This hierarchical architecture allows for the learning of multi-scale contextual representations, contributing to improved segmentation results. Despite its advantages, Swin UNETR also exhibits certain weaknesses. The model’s complexity and resource intensiveness pose challenges, requiring significant computational resources for training and inference due to the combination of Swin transformers and CNN-based decoders. Moreover, Swin UNETR’s performance heavily depends on the availability and quality of training data, making it susceptible to limitations or biases in datasets.
10. Dataset
| Ground Truth | LoGoNet | DiNTS Search | SwinUNETR48 | UNet++ | Attention U-Net |
This section is dedicated to providing a thorough exposition of the datasets meticulously chosen for inclusion in our experimental analyses. The selection of these datasets is pivotal in influencing the outcomes of our study, serving as the cornerstone for the evaluation and refinement of our methodologies. A detailed exploration of each dataset is undertaken to illuminate its characteristics and significance within the context of our experimental framework.
10.1. Unlabeled Datasets
As unlabeled data, we use a set of 4500 distinct data points re-published as a meta-dataset by Tang et al. (2022). Each of these data points represents a volumetric 3D scan specifically focusing on key anatomical areas such as the chest, abdomen, head, and neck. This deliberate selection of regions ensures a thorough examination of the anatomical aspects relevant to our research. This meta-dataset consists of the following dataset:
HNSCC 101010Available at: https://wiki.cancerimagingarchive.net/display/Public/HNSCC: imaging, radiation therapy, and clinical data of head and neck squamous cell carcinoma (HNSCC) patients at MD Anderson Cancer Center, the dataset comprises 954 samples.
LUNA16 111111Available at: https://luna16.grand-challenge.org/Data/: is a publicly available data set and challenge designed to advance the development of computer-aided detection (CAD) algorithms for the accurate identification of pulmonary nodules in low-dose computed tomography (CT) scans, contributing to the improvement of lung cancer detection, and the data set consists of 842 samples.
CT Colonography121212Available at: https://www.cancerimagingarchive.net/collections/: part of the National CT Colonography Trial, this dataset comprises 1532 cases of CT colonography imaging accompanied by polyp descriptions and their respective locations within colon segments. It serves as a valuable resource for validating the use of CT colonography in detecting colorectal neoplasia.
CT Images in COVID-19 131313TCIA COVID-19 Datasets: this collection of medical imaging data includes CT scans and associated clinical information collected during the COVID-19 pandemic. The dataset includes 722 samples.
LIDC-IDRI 141414Available at: wiki.cancerimagingarchive.net/display/Public/LIDC-IDRI: is a web-accessible collection of thoracic computed tomography (CT) scans with marked-up annotations of lung nodules, designed for the development and evaluation of computer-assisted diagnostic methods for lung cancer detection, the dataset comprises 450 samples.
10.2. BTCV Dataset
The provided dataset151515Available at: https://www.synapse.org/#!Synapse:syn3193805/wiki/217789 consists of 40 abdominal CT scans that were collected under Institutional Review Board (IRB) supervision. These scans are sourced from a combination of an ongoing colorectal cancer chemotherapy trial and a retrospective ventral hernia study. The CT scans were obtained during the portal venous contrast phase, capturing the anatomical details of the abdominal region. The dataset showcases variability in terms of volume sizes, with the scans ranging from to voxels. Additionally, the field of view (FOV) varies, spanning approximately to . In-plane resolution fluctuates between and , while slice thickness varies from to . Standard registration data for the dataset has been generated using NiftyReg.
For the preparation of the data set, we sliced at precise intervals of mm in both the x and y directions and mm for the isotropic resolution in the z-direction.
The dataset includes manual segmentations of thirteen abdominal organs. The segmentation was performed on a volumetric basis, providing accurate organ boundaries within the CT scans. The list of segmented organs includes the spleen, right kidney, left kidney, gallbladder, esophagus, liver, stomach, aorta, inferior vena cava, portal vein and splenic vein, pancreas, right adrenal gland, and left adrenal gland. It’s important to note that some patients within the dataset may not have a right kidney or gallbladder, and as a result, these organs might not be labeled in those specific cases.
The dataset serves as a valuable resource for advancing automated segmentation approaches in the field of medical imaging. Researchers can utilize this dataset to develop and refine algorithms that accurately identify and segment abdominal organs in CT scans.
| SSL Approach | Spl | RKid | Lkid | Gall | Eso | Liv | Sto | Aor | IVC | Veins | Pan | Rad | Lad | AVG |
[…]
| Rubik’s Cube (Tao et al., 2020) | .961 | .940 | .946 | .815 | .820 | .971 | .878 | .922 | .872 | .780 | .825 | .720 | .725 | .859 |
[…]
10.3. MSD Dataset
This dataset161616Available at: http://medicaldecathlon.com/ contains ten different tasks. Within this data set, a wide variety of organs and structures are known to encompass vital anatomical regions of significance, such as the heart, liver, prostate, and pancreas.
To use the dataset, we divided it into exact 1.0 mm intervals along the x, y, and z directions, maintaining isotropic resolution for all tasks.
We evaluated our models using six tasks from the MSD dataset. Table 15 displays the case numbers for the various tasks. The initial task focuses on ”Colon Cancer Primaries” and poses a challenge due to its varied and irregular appearance, referred to as ”heterogeneous” (heterogeneous indicating dissimilar components or elements, causing irregular or variegated appearances; for instance, a dermoid cyst exhibits heterogeneous attenuation on CT scans).
The subsequent task revolves around ”Spleen,” involving a challenge related to the significant variation in foreground size. The following task targets ”Hepatic vessels and tumor,” presenting a challenge in dealing with small tubular structures adjacent to the heterogeneous tumor. Moving on, we have the ”Pancreas Tumor” task, which aims to segment the liver and tumor, posing an unbalanced labeling challenge with large (background), medium (pancreas), and small (tumor) structures.
The ”Lung Tumors” task comes next, presenting the challenge of segmenting a small target (cancer) within a large image. Lastly, we have the ”Cardiac” task, which focuses on segmenting the Left Atrium and faces the challenge of a small training dataset with significant variability.
11. Qualitative and Quantitative Results
In this section, we report additionall qualitative and quantitative results of our model compared to the baseline models. These experiments we carried out in BTCV dataset. For all the experiments, to accommodate the constraints imposed by the limited RAM capacity of the GPUs employed for training, we implemented a cropping approach to diminish the size of the model. We opted for a cropping approach with dimensions of . To derive the final results, we employed a sliding window inference methodology, in which the required data were systematically passed with a mm overlapping mechanism.
In Tables 16, the results of our proposed model and pre-training approach are compared to the baselines. This analysis provides valuable insights into the effectiveness and performance gains achieved by our proposed methodology when compared against existing baseline methods.
The data presented in Table 17 illustrates the superiority of our innovative multi-task self-supervised learning strategy when compared to other baseline methods, as evidenced by the aggregated results on the BTCV dataset. Notably, our proposed approach showcases exceptional efficacy across three out of four experiments, highlighting its versatility and effectiveness across various tasks.
In Figure 5, the comprehensive results of our proposed model are presented alongside the outcomes of various baselines. Our model, equipped with a combined local and global attention mechanism, showcases superior performance in segmenting various body regions, including even the smaller anatomical parts. This enhanced segmentation capability is attributed to the ability of our model to effectively capture both local and long-range dependencies within the data, making it particularly proficient in delineating intricate structures within medical images. This mechanism provides a substantial advantage over the baselines, enabling our model to deliver more accurate and detailed segmentations across a wide range of body regions, which is essential for medical image analysis and diagnosis. The combination of global and local scales in our model, as well as its emphasis on meaningful feature extraction, makes it a powerful tool for improving the accuracy and precision of medical image segmentation tasks, particularly in scenarios where small anatomical details are crucial for accurate diagnosis.